A movie star, unwilling to give his age, posed the following riddle to a gossip columnist: "Eight years ago, I was fifteen times as old as my daughter. Now I am seven times as old as she is." How old is the movie star?
Answer by Boreal(15235)

movie star's age is x; daughter is y
x-8=15(y-8)
x=7y
Substitute into the first and distribute.
7y-8=15y-120
112=8y
y=14, daughter's age. Movie star would be 98. ANSWER.That is 1:7
8 years ago, daughter was 6, and movie star 90. That is 1:15.